A pure titanium cube has an edge length of 2.72 in .
How many titanium atoms does it contain? Titanium has a density of .4.50g/cm3
Answer #1

volume of Ti cube = 2.723 in3 = 20.12 in3 x 16.39 cm3/in3 = 329.8 cm3

density = 4.50 g/cm3

mass of Ti = density x volume = 4.50 x 329.8 = 1484 g

atomic mass of Ti = 47.88 g/mol

moles of Ti = mass / atomic mass = 1484/47.88 = 31 mol

1 mole of Ti contains 6.02 x 1023 Ti atoms,

Thus,

total Ti atoms number = 31 x 6.02 x 1023 = 1.87 x 1025
Answer #2
  • Once you convert the edge length (a) to centimetres, the volume in cubic centimetres is calculated as a3.

  • Multiplying the volume by the density of titanium yileds the mass of the cube in grams.

  • Then, the atomic mass of titanium can be used to convert the mass in grams to the number of moles.

  • Lastly, Avogadro's number is used to convert the number of moles to the number of titanium atoms.




The dimensional analysis for these calculations is summarised as follows:



(2.72 in • (2.54cm/1in))3 • (4.50g/1cm3) • (1 mol Ti/47.867g) • (6.022•1023 atoms/1mol)



Therfore, N = 1.87 • 1025 atoms Ti.

source: MyLab and Mastering: Chemistry
answered by: anonymouse